OpenStudy (anonymous):

Angle S and angle T are congruent. Angle S measures 2x + 5 and angle T measures 5x – 25. What is the measure of angle S?

OpenStudy (anonymous):

Step 1: What do we know? 1) Angles S and T are congruent. What does that mean? It means they are equal in measure, right? 2) We know Angle S measures 2x + 5 3) We know Angle T measures 5x - 25 Step 2: Use what we know... If angle S and T are congruent, then their measures should be equal. so using statements (2) and (3) above, 2x + 5 must be equal 5x - 25 2x+5 = 5x-25 Step 3: Solve for x. Step 4: Substitute value for x back into expression for Angle S = 2x + 5 to solve for measure of Angle S.

OpenStudy (anonymous):

i got -4.2 i know its not write i cant get it

OpenStudy (anonymous):

Is this a geometry class, or algebra, or what?

OpenStudy (anonymous):

geomertry

OpenStudy (anonymous):

So, surely sometime last year in algebra, you learned how to solve for x in an equation like: 2x + 5 = 5x - 25 You can subtract the 2x on the left from both sides of the equation... 2x - 2x + 5 = 5x - 2x - 25 5 = 3x - 25 Then add 25 to both sides of the equation to move the "25" from right to left... 5 + 25 = 3x - 25 + 25 30 = 3x Then divide by 3 to solve for x 30 / 3 = 3x / 3 10 = x Then you have to go back and solve for Angle S = 2x + 5 by using that x = 10. 2x + 5 = 2(10) + 5 = 25 Measure of Angle S = 25 degrees. If you can recall that basic algebra from last year, that's all you have to do here... algebra plus know that congruent angles have equal measure. If you don't remember the algebra, then you really need to refresh yourself on it... now's the time, because you will need it from now on.